The authors of a paper titled "Age and Violent Content Labels Make Video Games Forbidden Fruits for Youth" carried out an experiment to determine if restrictive labels on video games actually increased the attractiveness of the game for young game players.

Participants read a description of a new video game and were asked how much they wanted to play the game. The description also included an age rating. Some participants read the description with an age restrictive label of 7+, indicating that the game was not appropriate for children under the age of 7. Others read the same description, but with an age restrictive label of 12+, 16+, or 18+.

The summary data shown below are the ratings of 12- to 13-year-old boys for how much they wanted to play the game on a scale of 1 to 10. You can assume that boys were assigned at random to one of the four age label treatments (7+, 12+, 16+, 18+) and that ratings were normally distributed.

Group Sample Size Sample Mean Sample Std Dev
18+ label 22 7.2 1.33
16+ label 21 6.2 1.7
12+ label 18 6 1.5
7+ label 19 5 2.03

(f) Complete the ANOVA table. Note that you've already found two of the three sums of squares. Use three decimals for the sums of squares and means squares, then round the F test statistic to 2 decimals.

Source of Variation df Sum of Squares Mean Square F
Treatment
Error X
Total X X

(g) Do the data provide convincing evidence that the mean rating associated with the game description by 12- to 13-year-old boys is not the same for all four restrictive rating labels? Test the appropriate hypotheses using a significance level of 0.05.

test statistic:

Now we can test each mean against each other mean, and use a Bonferroni correction to control the family-wise error rate at α=0.05. Recall from above that:

(x)^^\_18+ = 7.2 (x)^^\_16+ = 6.2 (x)^^\_12+ = 6 (x)^^\_7+ = 5




For each comparison, calculate the difference in sample means, the t-test statistic (rounded to three decimal places), and the p-value. The standard error and degrees of freedom for each comparison are provided.

Parameter difference in sample means standard error of difference t-test statistic degrees of freedom two sided p-value
μ18+μ16+ 0.50 41
μ18+μ12+ 0.53 38
μ18+μ7+ 0.52 39
μ16+μ12+ 0.53 37
μ16+μ7+ 0.52 38
μ12+μ7+ 0.54 35
Answer #1


(f) SSE= sum of square due to error = (22 - 1)* 1.332 + (21 – 1) *1.72 + (18 - 1) * 1.52 + (19 - 1)* 2.032 = 207.373 G = rand

Source of Variation df Sum of Squares Mean Square F
Treatment 3 49.838 49.838/3=16.613 16.613/2.729=6.09
Error 76 207.373 207.373/76=2.729 X
Total 79 257.211 X X

(g) Critical value=F0.05,3,76=2.72

Since F=6.09>Critical value so the data provide convincing evidence that the mean rating associated with the game description by 12- to 13-year-old boys is not the same for all four restrictive rating labels.

Parameter difference in sample means standard error of difference t-test statistic degrees of freedom two sided p-value
μ18+μ16+ 7.2-6.2=1 0.50 1/0.5=2 41 0.0522
μ18+μ12+ 7.2-6=1.2 0.53 1.2/0.53=2.264 38 0.0294
μ18+μ7+ 7.2-5=2.2 0.52 2.2/0.52=4.231 39 0.0001
μ16+μ12+ 6.2-6=0.2 0.53 0.2/0.53=0.377 37 0.7083
μ16+μ7+ 6.2-5=1.2 0.52 1.2/0.52=2.308 38 0.0265
μ12+μ7+ 6-5=1 0.54 1/0.54=1.852 35 0.0725

Error Rate for each pair=0.05/6=0.0083

Hence μ18+ and μ7+ are significantly different (since p-value<0.0083) and other pairs are insignificantly different (since p-value>0.0083).